# Homogenized C. elegans Neural Activity and Connectivity Data

### v1

@@ 1 Introduction
Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience and machine learning. While neural dynamics in the brain can often be complex and difficult to measure in larger organisms, smaller systems like C. elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure [24, 3]. C. elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [25]. This makes it an ideal organism for studying how neural dynamics arise from connectivity and how this relationship could inform both neuroscience and AI models of neural function.
@@
With its transparent body and well-characterized neural architecture, C. elegans provides an excellent model for whole-brain functional imaging using calcium fluorescence sensors [26]. Calcium indicators act as proxies for voltage activity, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6]. Additionally, since the full connectome of C. elegans has been mapped via electron microscopy [24, 25], combining these two datasets—neural dynamics and connectomic structure—allows for a deep investigation of how network connectivity influences neural function. Unification of such datasets can also contribute to the broader goal of understanding how neural activity propagates through networks, ultimately informing both the development of AI-based neural systems and providing insights into how these mechanisms generalize to more complex organisms [8].
@@
However, the existing datasets from calcium imaging studies of C. elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates [9, 10]. This makes it challenging to perform cross-study comparisons or build unified models.
@@
To address these challenges, we have created two datasets, the first combining neural dynamics from 11 neuroimaging studies and the second combining neural connectivity from 10 connectome studies of C. elegans. This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data [5]. The primary goal of this dataset is to facilitate the training of foundation models of neural dynamics, which can be used to study how neural function arises from structure. The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel [15]. It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function [3]. We are releasing these datasets as open-source resources on the HuggingFace platform to ensure broad accessibility to researchers in both neuroscience and machine learning. By unifying these datasets and making them accessible, we hope to encourage the development of computational models that can capture the structure-function relationships in small nervous systems.
@@ 2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that supports the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions and files) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data in a standardized format.
@@
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The former handles the extraction and processing of the raw calcium data from each of our 12 neural activity source datasets into standardized neural activity data that are saved as compressed files. The latter processes the raw connectome data from connectome source files into compressed graph tensors that are saved as compressed in a data format amenable to downstream processing by graph neural network packages. There are 7 distinct connectome source files as shown in Table 5 but they are all derived from 3 primary source publications [24, 3, 25].
@@ 2.2.1 Neural Data Processing
The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$ enabling in its potential use in downstream analyses that require filtering out unlabeled neurons. Therefore, the pair $(\mathbf{M}^{(k)},\mathbf{X}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
@@ 2.3.1 Graph Tensor Data Processing
Table 5: Column Data Format for Connectome Dataset. Example of tabular data structure used in connectome datasets. Columns describe the origin and target neurons, connection type (gap junction or chemical synapse), number of connections, and associated neurotransmitters. The example shown is the first and last few rows of the OpenWorm [21]
connectome source file. Origin Target Type Number of Connections Neurotransmitter ADAL ADAR GapJunction 1 Generic_GJ ADAL ADFL GapJunction 1 Generic_GJ ADAL AIBL Send 1 Glutamate ADAL AIBR Send 2 Glutamate ADAL ASHL GapJunction 1 Generic_GJ VC1 VC2 Send 3 Serotonin_Acetylcholine VC1 VC2 GapJunction 6 Generic_GJ VC1 VC3 Send 1 Serotonin_Acetylcholine VC1 VC3 GapJunction 2 Generic_GJ VC1 VD1 GapJunction 1 Generic_GJ
@@ 3 Discussion
In this paper we present a normalized and standardized dataset for C. elegans neural activity and connectomes. We have created & consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code blocks we have provided are composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
This unified neural and connectome dataset provides a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. However, while this dataset offers numerous advantages, there are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. One notable clarification is that while C. elegans is often said to have 302 neurons, we refer to it as having 300 neurons in this work because two of these cells, CANL and CANR, are now classified as “end organs” rather than neurons. This shift in classification is supported by the work of Stefanakis et al. [17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
@@
Additionally, while the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and potentially across the lifespan of a single individual. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [3, 14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity. This limitation means that the relationship between structure and function is inferred across different animals, potentially missing individual-specific relationships between connectivity and function, despite phenotypic matching. The generation of the C. elegans connectome, as described by White et al. [24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons due to neuromodulator & neuropeptide release which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [14].

### v2

@@ 1 Introduction
Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience and machine learning. While neural dynamics in the brain can often be complex and difficult to measure in larger organisms, smaller systems like C. elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure. C. elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [24, 3, 25]. This makes it an ideal organism for studying how neural dynamics arise from connectivity and how this relationship could inform both neuroscience and AI models of neural function.
@@
With its transparent body and well-characterized neural architecture, C. elegans provides an excellent model for whole-brain functional imaging using calcium fluorescence sensors [26]. Calcium indicators act as proxies for local computation, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6]. Additionally, since the full connectome of C. elegans has been mapped via electron microscopy [24, 3, 25], combining these two datasets — neural dynamics and connectivity structure — allows for a deep investigation of how network connectivity influences neural function. Unification of such datasets can also contribute to the broader goal of understanding how neural activity propagates through networks, ultimately informing both the development of AI-based neural systems and providing insights into how these mechanisms generalize to more complex organisms [8].
@@
However, the existing datasets from calcium imaging studies of C. elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates. This makes it challenging to perform cross-study comparisons or build unified models.
@@
To address these challenges, we have created two datasets, the first combining neural dynamics from 11 calcium maging studies and the second combining neural connectivity from 3 connectome studies of C. elegans [24, 3, 25]. This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data. The primary goal of this dataset is to facilitate the training of computational models of C. elegans neural dynamics, which can be used to study how neural function is related to structure. The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel. It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function. We are releasing these datasets as open-source resources on the HuggingFace platform to ensure broad accessibility to researchers in both neuroscience and machine learning. By unifying these datasets and making them accessible, we hope to encourage the development of computational models that can capture the structure-function relationships in small nervous systems.
@@ 2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions and files) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data in a standardized format.
@@
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The former handles the extraction and processing of the raw calcium data from each of our 11 neural activity source datasets into standardized neural activity data that are saved as compressed files. The latter processes the raw connectome data from connectome source files into compressed graph tensors that are saved as compressed in a data format amenable to downstream processing by graph neural network packages. There are 10 distinct connectome source files as shown in Table 3 but they are all derived from 4 primary source publications [24, 3, 25, 14].
@@ 2.2.1 Neural Data Processing
The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$ enabling in its potential use in downstream analyses that require filtering out unlabeled neurons. Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
@@ 2.3.1 Graph Tensor Data Processing
Table 5: Column Data Format for Connectome Dataset. Example of tabular data structure used in connectome datasets. Columns describe the origin and target neurons, connection type (gap junction or chemical synapse), number of connections, and associated neurotransmitters. The example shown is the first and last few rows of the OpenWorm [21] connectome source file
. Origin Target Type Number of Connections Neurotransmitter ADAL ADAR GapJunction 1 Generic_GJ ADAL ADFL GapJunction 1 Generic_GJ ADAL AIBL Send 1 Glutamate ADAL AIBR Send 2 Glutamate ADAL ASHL GapJunction 1 Generic_GJ VC1 VC2 Send 3 Serotonin_Acetylcholine VC1 VC2 GapJunction 6 Generic_GJ VC1 VC3 Send 1 Serotonin_Acetylcholine VC1 VC3 GapJunction 2 Generic_GJ VC1 VD1 GapJunction 1 Generic_GJ
@@ 3 Discussion
In this paper we present a normalized and standardized dataset for C. elegans neural activity and connectomes. We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code blocks we have provided are composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
This unified neural and connectome dataset provides a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. However, while this dataset offers numerous advantages, there are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. One notable clarification is that while C. elegans is often said to have 302 neurons, we refer to it as having 300 neurons in this work because two of these cells, CANL and CANR, are now classified as “end organs” rather than neurons. This shift in classification is supported by [17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
@@
Additionally, while the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and potentially across the lifespan of a single individual. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [3, 14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity. This limitation means that the relationship between structure and function is inferred across different animals, potentially missing individual-specific relationships between connectivity and function, despite phenotypic matching. The generation of the C. elegans connectome, as described by [24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons due to neuromodulator & neuropeptide release which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [14].